# Diagnosing and Mitigating Modality Interference in Multimodal Large Language Models

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) have demonstrated impressive capabilities across tasks, yet they often exhibit difficulty in distinguishing task-relevant from irrelevant signals—particularly in tasks like Visual Question Answering (VQA)—which can lead to susceptibility to misleading or spurious inputs. We refer to this broader limitation as the Cross-Modality Competency Problem—the model’s inability to fairly evaluate all modalities. This vulnerability becomes more evident in modality-specific tasks—such as image classification or pure text question answering—where models are expected to rely solely on one modality. In such tasks, spurious information from irrelevant modalities often lead to significant performance degradation. We refer to this failure as Modality Interference, which serves as a concrete and measurable instance of the cross-modality competency problem, and we further design a perturbation-based causal diagnostic experiment to verify and quantify this problem. To mitigate modality interference, we propose a novel framework to finetune MLLMs, including perturbation-based data augmentations with both heuristic perturbations and adversarial perturbations, and a consistency regularization strategy applying on model outputs with original and perturbed inputs. Experiments on multiple benchmark datasets (image-heavy, text-heavy and multimodal tasks) and multiple model families with different scales demonstrate significant improvements in robustness and cross-modality competency, indicating our method’s effectiveness in boosting unimodal reasoning ability while enhancing performance on multimodal tasks. Our code and models can be found at: https://github.com/luisrui/Modality-Interference-in-MLLMs
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have made significant strides in integrating vision and language understanding within a unified architecture  (Liu et al., 2023b; Luo et al., 2023; Bai et al., 2025). By combining powerful visual encoders and large language models through alignment mechanisms, MLLMs such as LLaVA (Liu et al., 2023b) and Qwen-VL (Bai et al., 2025) demonstrate strong capabilities across a wide range of multimodal tasks. However, beneath their seemingly impressive performance lies a critical limitation: MLLMs often fail to distinguish between relevant and irrelevant signals across modalities, leading to unreliable predictions (Wang et al., 2024a; Zhu et al., 2024; Hosseini et al., 2025). Moreover, while MLLMs are designed for multimodal tasks, their failure on unimodal tasks—where only a single modality (e.g. text) should guide the prediction—raises concerns about whether the model can preserve modality-specific competencies. For instance, MLLMs frequently underperform on pure visual recognition (Zhang et al., 2024; Tong et al., 2024) and textual reasoning (Zhu et al., 2024; Wang et al., 2023; Lin et al., 2024), suggesting that cross-modal fusion may induce unintended interference and degrade unimodal performance.
@@
Recent studies have attributed this phenomenon to a variety of symptoms arising during the vision-language alignment process, such as catastrophic forgetting (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023; Lin et al., 2024), knowledge conflict (Wang et al., 2024a; Zhu et al., 2024), and spurious correlations (Chen et al., 2024a; Hosseini et al., 2025). Catastrophic forgetting has been identified as a key factor in visual degradation (Zhang et al., 2024; Tong et al., 2024), where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict (Wang et al., 2024a; Zhu et al., 2024) impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations (Chen et al., 2024a; Hosseini et al., 2025; Zhou et al., 2025) show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding. While these symptoms shed light on MLLMs’ limitations, most works treat these issues in isolation. For instance, architectural issues such as shallow cross-modal fusion have been widely discussed: lightweight projectors in models like LLaVA (Liu et al., 2023b) fail to fully align vision and language representations, resulting in unstable modality reliance (Tong et al., 2024; Zhu et al., 2024; Zhao et al., 2025). Others attribute performance bottlenecks to data limitations—insufficient modality-specific supervision leads to impaired visual decoding (Zhang et al., 2024) and diminished language understanding (Lin et al., 2024). Inspired by these observations, our insight is to unify these challenges under a broader perspective: the model’s inability to identify and rely on the modality that contributes most relevant information to the task. We argue that the fundamental limitation lies in MLLMs’ lack of cross-modality competency (Gardner et al., 2021)—the ability to fairly evaluate and integrate information across modalities. Current MLLMs lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
Figure 1: Performance degradation under irrelevant perturbations reveals modality interference in MLLMs. Left: Mini-ImageNet (image-heavy) with Original input, Unrelated Facts, and Misleading Descriptions. Right: OpenBookQA (text-heavy) with Random Pixels, Full Black Canvas, and Irrelevant Real Images. Misleading descriptions induces the most severe degradation in image-heavy tasks, while irrelevant real images cause the largest drop in text-heavy reasoning.
@@
To systematically diagnose and mitigate modality interference, we introduce a two-stage methodology grounded in causal analysis. First, we design a perturbation-based evaluation experiment inspired by causal intervention principles (Pearl, 1995; Chen et al., 2024a) to diagnose the extent of modality interference across tasks and model scales. Second, we propose a robust fine-tuning framework to mitigate modality interference. Specifically, in our evaluation analysis, we first focus on modality-heavy settings using a multiple-choice question answering format, where the model selects an answer from a fixed set of options based on both image and text input. We then include image-heavy tasks (e.g., image classification), text-heavy tasks (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), allowing us to examine how models behave under different modality-reliance scenarios. To further induce modality interference, we introduce heuristic perturbations: In image-heavy tasks, we perturb the text input by prepending either (i) unrelated scientific facts or (ii) misleading descriptions that falsely associate an incorrect option with the image content. In text-heavy tasks, where the default visual input is random noise, we perturb the visual input with (i) semantically meaningful real images, (ii) full black canvas, or (iii) full white canvas. These perturbations are designed to either introduce spurious cues or reinforce irrelevant modality signals. We evaluate the resulting changes in model predictions to assess the robustness of modality selectivity. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model modality interference through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs across different architectures and scales with results shown in Figure 1. In image-heavy tasks, unrelated textual facts moderately reduce performance, while misleading descriptions cause severe degradation—revealing the model’s vulnerability to spurious textual cues. In text-heavy tasks, canvas inputs have little effect, but unrelated real images mostly hurt performance, indicating improper fusion of irrelevant visual signals into reasoning.
@@
The empirical results from Figure 1 confirm the presence of modality interference and reveal the limitations of current MLLMs in lack of cross-modality competency. To mitigate modality interference, we propose a perturbation-based fine-tuning framework for MLLMs. Specifically, to alleviate data insufficiency, we apply a perturbation-based data augmentation strategy, where we construct a diverse mixture of modality-specific, perturbation-augmented samples and original VQA samples. The perturbations include both heuristic variants (e.g., injecting unrelated facts into image-heavy prompts) and adversarial training-time perturbations, which expose the model to worst-case alignment disruptions and thus serve as a stronger form of regularization. To further improve robustness, we introduce a consistency regularization strategy(e.g., via Jensen–Shannon divergence), which enforces output stability between original and perturbed samples. In summary, the main contributions of this paper are threefold. First, we introduce the notion of the Cross-Modality Competency Problem to describe how multimodal models may struggle to balance different modalities, and analyze modality interference as one concrete instance of such challenges in MLLMs. Second, we design a perturbation-based causal evaluation experiment that systematically quantifies modality reliance and reveals models’ susceptibility to modality interference. Third, we propose a fine-tuning strategy that combines supervised augmentation with both heuristic and adversarial perturbations and consistency regularization to mitigate modality interference. Extensive experiments across multiple MLLM families and diverse benchmarks demonstrate the superiority of our method.
@@ 2 Related Works
Recent studies have revealed that modality misalignment remains a key obstacle in MLLMs, leading to degraded performance on both image-heavy and text-heavy tasks. For visual understanding, catastrophic forgetting occurs when multimodal tuning overrides pretrained visual features (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023). In text-heavy scenarios, knowledge conflict (Zhu et al., 2024) arises when inconsistent parametric knowledge from different modalities confuse reasoning. Moreover, mDPO (Wang et al., 2024a) identifies language bias in training, where models fail to condition their responses on visual input. Some works attribute such issues to shallow fusion (Wang et al., 2023)—e.g., LLaVA uses lightweight projectors to bridge vision and language spaces, leaving a representational gap and resulting in loosely coupled features (Tong et al., 2024; Zhu et al., 2024; Zhao et al., 2025). Others highlight data limitations: even well-encoded visual features fail to support reasoning without adequate supervision to guide decoding (Zhang et al., 2024).
Building on these diagnoses, recent models have proposed multiple solutions. MoF (Tong et al., 2024) mitigates this by fusing features from multiple vision encoders, while VLMClassifier (Zhang et al., 2024) enhances recognition via vision-only finetuning, though it struggles with VQA due to lack of cross-modal alignment. CogVLM (Wang et al., 2023) introduces a visual expert module to improve vision-language integration. VILA (Lin et al., 2024), QwenVL (Bai et al., 2025), and InternVL (Zhu et al., 2025) incorporate text-only supervision in different ways to preserve or enhance language capabilities during multimodal training—through stage-wise separation, parallel preservation, and unified joint optimization, respectively. Similar patterns arise in multimodal structural reasoning, where models must rely on relevant modalities to generalize to unseen relations (Cai et al., 2024). These works motivate us to hypothesize on the root cause—the model’s inability to assess modality relevance. We further propose a causal framing of modality interference and introduce a perturbation-based fine-tuning strategy to improve the inference-time robustness of MLLMs.
@@
Adversarial perturbations threaten the reliability of both vision and text tasks by exposing vulnerabilities through deliberate and imperceptible perturbations. Existing robustness methods can be broadly categorized by modality, targeting either continuous image embeddings or discrete token spaces. In vision tasks, attacks like FGSM (Goodfellow et al., 2015) and CW (Carlini & Wagner, 2017) first revealed the fragility of neural networks to imperceptible input changes. PGD (Madry et al., 2018) formalized this under a saddle-point framework, becoming the standard for adversarial training. AutoAttack (Croce & Hein, 2020) further unified strong attacks, including PGD variants, into a reliable benchmark. In text tasks, adversarial methods must contend with discrete inputs. TextFooler (Jin et al., 2020) substitutes key words with semantically similar ones to mislead predictions, while CodeAttack (Jha & Reddy, 2023) adapts this idea to code-language models. More recently, PGD has been extended to LLMs via continuous relaxation (Geisler et al., 2024), enabling efficient attacks in embedding space. Beyond evaluation, PGD has also been used as a regularizer to improve optimization. PTP (Chen et al., 2023a) applies PGD-style perturbations in the prompt embedding space to smooth training and enhance stability. Inspired by this, our work extends PGD to the multimodal embedding space, enabling unified gradient-based control over both visual and textual inputs.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features (Gardner et al., 2021). We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency as an ability for MLLM to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. For instance, in a pure-text question answering task, the model receives both a question and an image, as is standard in MLLM input formats. However, the image is not required to answer the question. If the model relies on spurious visual cues—such as objects or scenes that frequently co-occur with certain answers—it violates the task’s competency condition by grounding predictions in irrelevant modality signals. This manifests as Modality Interference, where the presence of an irrelevant but misleading modality disrupts the model’s reasoning.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. Details in Section 5, Section C.1 and Table 8.
@@ 4.2 Consistency Regularization under Perturbations
$\mathcal{L}_{\text{consistency}}=\frac{1}{L_{A}}\sum_{i=1}^{L_{A}}\sum_{v=1}^{V}\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}\cdot\log\frac{\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}}{\text{softmax}\!\left(\tfrac{l^{\text{pert}}_{i}}{\tau}\right)_{v}}.$ (8)
4.3 Final Training Objective
Our final training objective integrates both perturbation-based data augmentation and consistency regularization into a unified framework. For each batch, we begin with a set of original samples $\mathcal{B}_{\text{orig}}$ and dynamically construct their heuristic perturbed counterparts $\mathcal{B}_{\text{pert}}$ via input-level augmentations. We then apply adversarial perturbations on both $\mathcal{B}_{\text{orig}}$ and $\mathcal{B}_{\text{pert}}$, and enforce consistency between the predictions of original and all perturbed samples. The overall loss is:
Table 1: Evaluation on unimodal and VQA datasets. For unimodal datasets, we report accuracy under the original input (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on the original setting. Best results are highlighted in bold.
@@ 4.3 Final Training Objective
5 Experiments
Models. We conduct experiments on three MLLM families with different parameter size: Qwen2.5-vl-3b (Bai et al., 2025), LLaVA-1.5-7B & LLaVA-1.5-13B (Liu et al., 2023a) and InstructBLIP-Vicuna-7B (Luo et al., 2023). Following LLaVA and Qwen-VL, we freeze the vision encoder and train the multimodal projector and language model; for InstructBLIP, we instead freeze both the vision encoder and Q-Former, fine-tuning only the projection layer and language model (see Appendix B).
Baselines. We include following baselines for comparison: LLaVA-1.5-13B + I-MoF (Tong et al., 2024): By applying the designed Interleaved Mixture-of-Features (I-MoF) module on LLaVA-1.5-13B to spatially combine CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023) visual tokens, it enhances visual grounding by integrating complementary features from contrastive and self-supervised vision encoders. VLMClassifier (Zhang et al., 2024): it enhances visually-grounded language models for image classification by fine-tuning them on ImageNet (Deng et al., 2009) (VLMClassifier-1) or ImageNet combining LLaVA-Instruct (Liu et al., 2023b) (VLMClassifier-2). Chain-of-Thought (CoT) Prompting (Wei et al., 2022): we further evaluate prompt-based mitigation by encouraging structured reasoning through CoT-style prompting. The specific prompt design and results are reported in Section C.4.
Datasets. We evaluate models on benchmarks covering three task types: (i) Image-heavy tasks: Mini-ImageNet (Russakovsky et al., 2015) and Caltech-101 (Fei-Fei et al., 2004), used in both training and evaluation, originally designed for image classification; (ii) Text-heavy tasks: OpenBookQA (Mihaylov et al., 2018) and MMLU (Hendrycks et al., 2020), consisting purely of textual question answering data; (iii) VQA tasks: For training, we use LLaVA-Instruct-dataset (Liu et al., 2023b) as the instruction-tuning dataset for related models. For InstructBLIP, we additionally use TextCaps (Sidorov et al., 2020) as another publicly available VQA dataset used in instruction tuning. For Qwen2.5-VL, whose instruction-tuning data is proprietary, we adopt LLaVA-Instruct as a standardized alternative. For evaluation, we adopt three multiple-choice VQA benchmarks: ScienceQA-IMG (Lu et al., 2022), MM-Bench-EN (Liu et al., 2023c), and Seed-Bench-IMG (Li et al., 2023). For ScienceQA and Seed-Bench, we only include examples with image context. For MM-Bench, we use the English version. All datasets are converted into a unified multiple-choice VQA format, enabling consistent modeling and evaluation across tasks and models. We report the accuracy of all multiple choice tasks and quantify the causal effect with the prediction change rate $\delta_{\text{cp}}$. All models are fine-tuned for 1 epoch with a fixed batch size $N_{\text{batch}}$. All the hyperparameters are listed in Appendix D. For each dataset, results are averaged over multiple independent runs. Following standard practice, inference is performed with deterministic decoding (temperature fixed at 0).
Achieving Pareto-Optimality Across Unimodal and Multimodal Tasks
As shown in Table 1, our method outperforms all baselines across different base MLLMs, demonstrating stronger robustness to modality interference and improved cross-modality competency. While CoT slightly improves performance in certain VQA settings, its overall gains are minimal and inconsistent, and it fails to mitigate modality interference under perturbed conditions (e.g., 85.9% vs. 97.9% on Mini-ImageNet). While I-MoF enhances visual grounding by integrating multiple visual features, it still suffers from modality interference: e.g. LLaVA-1.5-13B + I-MoF achieves 93.9% on original Mini-ImageNet but drops to 70.1% under perturbation (↓23.8%), indicating reliance on spurious textual cues. In contrast, our method maintains perturbed performance at 98.4% (↓0.1%). On the other hand, VLMClassifier, adopts vision-only fine-tuning, which leads to two critical limitations: vulnerability to cross-modal interference and degradation on VQA tasks, as LLaVA-1.5-7b + VLMClassifier-1 only reaches 35.8%/36.2% on MM-Bench/SeedBench, notably lower than both base LLaVA and our method (73.7%/68.4%). These results highlight that vision-centric strategies, without addressing modality alignment, are insufficient for robust multimodal understanding. In text-heavy tasks such as OpenBookQA and MMLU, our method also achieves superior perturbed performance(e.g., 55.8% vs. 39.1% on MMLU on LLaVA-1.5-13B)—highlighting that addressing modality interference directly, rather than merely improving representations, is key to robust multimodal reasoning. Overall, unlike prior methods that often trade off between unimodal and multimodal performance, our method consistently improves both, achieving Pareto-optimality.

### v4

@@ Abstract
Multimodal Large Language Models demonstrate strong performance on multimodal benchmarks, yet often exhibit poor robustness when exposed to spurious modality interference, such as irrelevant text in vision understanding, or irrelevant visual content in question answering. At its core, modality interference refers to cases where spurious signals from non-essential modalities distort model decisions, which we systematically analyze through causal, perturbation-based diagnostic experiments. To address this problem, we propose a unified finetuning framework that combines heuristic and adversarial perturbation-based data augmentation with output-level consistency regularization between original and perturbed inputs. Extensive experiments across image-heavy, text-heavy, and multimodal benchmarks, spanning multiple MLLM architectures and model scales, demonstrate consistent improvements in unimodal robustness and generalization, while improving standard multimodal performance.
Multimodal Large Language Models, Modality Interference, Robustness
@@ 1 Introduction
Multimodal Large Language Models (MLLMs) have made significant strides in integrating vision and language understanding within a unified architecture  (Liu et al., 2023b; Luo et al., 2023; Bai et al., 2025b). By combining powerful visual encoders and large language models through alignment mechanisms, MLLMs such as LLaVA (Liu et al., 2023b) and Qwen-VL (Bai et al., 2025b) demonstrate strong capabilities across a wide range of multimodal tasks. However, beneath their seemingly impressive performance lies a critical limitation: MLLMs often fail to distinguish between task-relevant and task-irrelevant signals across modalities, leading to unreliable predictions (Wang et al., 2024a; Zhu et al., 2024; Hosseini et al., 2025). A concrete manifestation of this limitation is their behavior in unimodal diagnostic settings, where only a single modality is required for correct reasoning. In such settings, MLLMs frequently underperform on pure visual recognition (Zhang et al., 2024; Tong et al., 2024) and textual reasoning (Zhu et al., 2024; Wang et al., 2023; Lin et al., 2024), indicating that non-essential modalities can still influence the prediction process.
@@
Recent studies have attributed this failure mode to a variety of learning challenges arising during the multimodal alignment process, such as catastrophic forgetting of pretraining knowledge (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023; Lin et al., 2024), cross-modality knowledge conflict (Wang et al., 2024a; Zhu et al., 2024), and spurious correlations (Chen et al., 2024a; Hosseini et al., 2025). Catastrophic forgetting has been identified as a key factor in visual degradation (Zhang et al., 2024; Tong et al., 2024), where multimodal tuning of MLLM overrides its pretrained visual features. Cross-modal knowledge conflict (Wang et al., 2024a; Zhu et al., 2024) impairs pure-text reasoning, as models often produce inconsistent outputs when visual inputs are introduced, reflecting misaligned visual and textual parametric memories. Additionally, studies on spurious correlations (Chen et al., 2024a; Hosseini et al., 2025; Zhou et al., 2025) show that MLLMs tend to rely on superficial cross-modal cues rather than task-relevant grounding.
Despite these extensive efforts, existing explanations largely remain fragmented, attributing the observed failures to isolated challenges arising during multimodal alignment. As a result, it remains unclear whether these seemingly diverse issues share a common inference-time mechanism that governs how MLLMs identify and rely on task-relevant modalities during prediction. Motivated by this gap, we adopt a unified perspective and argue that the core limitation lies in MLLMs’ lack of cross-modality competency (Gardner et al., 2021)—the ability to fairly evaluate and integrate information across modalities. Current models lack mechanisms to support this competency during inference, making them vulnerable to misleading cross-modal signals—a failure mode we refer to as Modality Interference.
Figure 1: Modality Interference in MLLMs. We visualize the performance of 15 MLLMs using radar charts, where the polygon area signifies model capability. Left (Vision Tasks): When vision tasks are interfered with by Misleading Text (red), models exhibit a severe performance collapse, shrinking towards the center. Right (Text Tasks): Conversely, text reasoning tasks suffer from Visual Noise (brown/irrelevant image), causing a noticeable degradation compared to the no-interference baseline.
@@
To systematically diagnose and mitigate modality interference, we first design a perturbation-based evaluation experiment inspired by causal intervention principles (Pearl, 1995; Chen et al., 2024a), covering multiple tasks and model scales. Specifically, in our analysis, we focus on modality-heavy settings using multiple-choice question answering across image-heavy (e.g., image classification), text-heavy (e.g., pure-text QA), and balanced multimodal tasks (e.g., VQA), enabling controlled analysis under different modality-reliance scenarios. To further induce modality interference, we inject heuristic perturbations (e.g., misleading textual descriptions or distractor visual inputs) into task-irrelevant modalities to introduce spurious or misleading signals, and quantify their impact on model predictions. We evaluate the resulting changes in model predictions to assess the extent of modality interference. While the perturbation-based evaluations offer empirical insights, we further frame our analysis through a causal intervention framework and in which we model this problem through a causal graph abstraction. Building on this framework, we evaluate a range of pretrained MLLMs and observe consistent failure patterns across tasks and scales, as shown in Figure 1. As illustrated, current MLLMs exhibit significant vulnerability to modality interference: misleading textual descriptions trigger a performance collapse in image-heavy tasks, while irrelevant visual noise disrupts text-heavy reasoning.
@@
The empirical results from Figure 1 confirm the presence of modality interference due to their lack of cross-modality competency. To mitigate this, we propose a perturbation-robust fine-tuning framework. We construct a diverse training mixture incorporating both heuristic noise and adversarial worst-case disruptions to simulate alignment failures, while simultaneously employing consistency regularization (e.g., via Jensen–Shannon divergence) to enforce output stability between clean and perturbed inputs. In summary, the main contributions of this paper are threefold. First, we introduce the Cross-Modality Competency Problem to characterize MLLMs’ difficulty in appropriately integrating information across modalities, and identify modality interference as a measurable manifestation of this issue. Second, we propose a causal, perturbation-based evaluation experiment that exposes inappropriate modality reliance and systematically quantifies MLLMs’ susceptibility to task-irrelevant signals. Third, we develop a dataset-agnostic, perturbation-agnostic fine-tuning strategy to mitigate modality interference. Extensive experiments across multiple MLLM families and diverse benchmarks consistently demonstrate the effectiveness and generality of our approach.
@@ 2 Related Works
Recent studies have revealed that modality misalignment remains a key obstacle in MLLMs, leading to degraded performance on both image-heavy and text-heavy tasks. For visual understanding, catastrophic forgetting occurs when multimodal tuning overrides pretrained visual features (Zhang et al., 2024; Tong et al., 2024; Wang et al., 2023). In text-heavy scenarios, knowledge conflict (Zhu et al., 2024) arises when inconsistent parametric knowledge from different modalities confuse reasoning. mDPO (Wang et al., 2024a) identifies language bias in training, where models fail to condition their responses on visual input. Some works attribute such issues to shallow fusion (Wang et al., 2023)—e.g., LLaVA (Liu et al., 2023b) uses lightweight projectors to bridge vision and language spaces, leaving a representational gap and resulting in loosely coupled features (Tong et al., 2024; Zhu et al., 2024; Zhao et al., 2025). Others highlight data limitations: even well-encoded visual features fail to support reasoning without adequate supervision to guide decoding (Zhang et al., 2024). Building on these diagnoses, recent works propose multiple solutions. MoF (Tong et al., 2024) mitigates this by fusing features from multiple vision encoders, while VLMClassifier (Zhang et al., 2024) enhances recognition via vision-only finetuning, though it struggles with VQA due to lack of cross-modal alignment. CogVLM (Wang et al., 2023) introduces a visual expert module to improve vision-language integration. VILA (Lin et al., 2024), QwenVL (Bai et al., 2025b), and InternVL (Zhu et al., 2025) incorporate text-only supervision in different ways to preserve or enhance language capabilities during multimodal training—through stage-wise separation, parallel preservation, and unified joint optimization, respectively. Similar patterns also arise in multimodal structural reasoning (Cai et al., 2024). However, these approaches largely treat modality contributions implicitly, without explicitly modeling how multimodal inputs influence the model’s decision process.
@@
Adversarial perturbations threaten the reliability of both vision and text models by exploiting vulnerabilities in continuous image embeddings and discrete token spaces. In vision tasks, attacks like FGSM (Goodfellow et al., 2015) and CW (Carlini and Wagner, 2017) first revealed the fragility of neural networks to imperceptible input changes. PGD (Madry et al., 2018) formalized this under a saddle-point framework, becoming the standard for adversarial training. AutoAttack (Croce and Hein, 2020) further unified strong attacks, including PGD variants, into a reliable benchmark. In text tasks, adversarial methods must contend with discrete inputs. TextFooler (Jin et al., 2020) substitutes key words with semantically similar ones to mislead predictions, while CodeAttack (Jha and Reddy, 2023) adapts this idea to code-language models. More recently, PGD has been extended to LLMs via continuous relaxation (Geisler et al., 2024), enabling efficient attacks in embedding space. PTP (Chen et al., 2023a) applies PGD-style perturbations in the prompt embedding space to smooth training and enhance stability. Inspired by this, our work extends to the multimodal embedding space, enabling unified gradient-based control over both visual and textual inputs with unique masking strategy.
@@ 3 Causal Analysis on Modality Interference
Competency problems describe scenarios where models rely on spurious correlations between isolated input features and output labels to make predictions, instead of leveraging meaningful interactions among multiple features (Gardner et al., 2021). We extend this concept to the multimodal setting by treating entire modalities (e.g., image $X_{I}$ or text $X_{T}$) as structured feature sources. We define the Cross-Modality Competency as an ability for MLLM to fairly evaluate and integrate all modalities, identifying which ones carry task-relevant signals while ignoring misleading or irrelevant ones. In pure-text question answering with MLLMs, the model is provided with both a question and an image, although the image is not required to solve the task. When predictions are influenced by spurious visual cues correlated with certain answers, the model allows irrelevant modality signals to influence its reasoning and predictions, resulting in modality interference.
@@
To systematically measure cross-modality competency, we propose a perturbation-based evaluation framework. The core idea is to inject controlled noise into the irrelevant modality and assess the model’s robustness to such perturbations. Specifically, for image-heavy tasks, we perturb the text input by: (1)Prepending unrelated scientific facts; (2) Prepending misleading descriptions that falsely link incorrect options to the image content. For text-heavy tasks, we perturb the visual input by: (1) Attaching a semantically meaningful real but irrelevant image; (2) Substituting with a full black or full white canvas image. Models with strong modality selectivity should maintain high prediction consistency when irrelevant modality signals are perturbed. We select multiple image-heavy and text-heavy tasks for evaluation. Each task is framed as a multiple-choice classification problem, requiring the model to choose the correct option based on image and text modalities as input, with perturbations applied as described above. To validate robustness and generalization for our proposed mitigation method, we additionally consider real-world out-of-distribution perturbations reflecting practical noise encountered in deployed multimodal systems: (i) noisy OCR snippets for image-heavy tasks; and (ii) unrelated UI screenshots for text-heavy tasks. Details in Section 5, Section C.1, Table 8 and Table 11 and Table 12.
@@ 4.2 Consistency Regularization under Perturbations
$\small\mathcal{L}_{\text{consistency}}=\frac{1}{L_{A}}\sum_{i=1}^{L_{A}}\sum_{v=1}^{V}\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}\cdot\log\frac{\text{softmax}\!\left(\tfrac{l^{\text{orig}}_{i}}{\tau}\right)_{v}}{\text{softmax}\!\left(\tfrac{l^{\text{pert}}_{i}}{\tau}\right)_{v}}.$ (8)
Table 1: Evaluation on unimodal and VQA datasets. For unimodal datasets, we report accuracy under the original input (Orig) and the worst-performing perturbation (Perturbed). For VQA datasets, we report accuracy on original setting. Best results are highlighted in bold.
@@
Qwen2.5-VL-7B 99.3 96.3 99.1 97.2 85.9 77.5 69.3 63.7 85.1 86.6 77.0
   + CoT (Wei et al., 2022) 99.1 91.4 98.5 92.5 77.8 68.8 61.5 57.4 85.2 86.7 77.1
R1-OneVision-7B (Yang et al., 2025) 98.2 87.7 96.5 84.7 81.2 77.6 63.2 60.8 84.5 85.4 76.3
Ours 99.6 99.5 99.6 99.7 90.9 89.7 69.8 68.2 83.9 87.0 78.3
4.3 Final Training Objective
Our final training objective integrates both perturbation-based data augmentation and consistency regularization into a unified framework. For each batch, we begin with a set of original samples $\mathcal{B}_{\text{orig}}$ and dynamically construct their heuristic perturbed counterparts $\mathcal{B}_{\text{pert}}$ via input-level augmentations. We then apply adversarial perturbations on both $\mathcal{B}_{\text{orig}}$ and $\mathcal{B}_{\text{pert}}$, and enforce consistency between the original and all perturbed samples. The overall loss is: